# Object-Centric Latent Action Learning

### v2

@@ 4 Experiments
Table 1: Normalized evaluation returns (DCS) and normalized success rates (DMW) of the BC agent trained on latent actions for different tasks. TL;DR: LAPO’s performance drops in the presence of distractors. Object-centric learning, using slots and masks, reduces this gap. Each value represents the average score across different amounts of fine-tuning labeled trajectories and three random seeds. The BC agent trained with access to the full dataset of ground-truth actions would achieve a score of 1 for each task. The relative performance gain is given as ${\frac{\text{LAPO}\ -\ \text{LAPO-slots/masks}}{\text{LAPO}\ -\ \text{LAPO-vanilla}}}$.
Task LAPO LAPO-vanilla LAPO-masks LAPO-slots baseline upper-bound ours ours cheetah-run 0.24 ± 0.02 0.76 ± 0.04 0.41 ± 0.03 +32% 0.55 ± 0.04 +58% hopper-hop 0.03 ± 0.01 0.27 ± 0.03 0.08 ± 0.01 +20% 0.15 ± 0.02 +50% walker-run 0.04 ± 0.01 0.32 ± 0.07 0.06 ± 0.01 +6% 0.12 ± 0.01 +27% humanoid-walk 0.02 ± 0.01 0.06 ± 0.01 0.04 ± 0.02 +47% 0.06 ± 0.01 +105% Mean DCS-hard 0.08 ± 0.01 0.35 ± 0.04 0.15 ± 0.02 +26% 0.22 ± 0.02 +52% cheetah-run 0.34 ± 0.06 0.76 ± 0.04 0.54 ± 0.09 +48% 0.50 ± 0.07 +38% hopper-hop 0.05 ± 0.01 0.27 ± 0.03 0.15 ± 0.01 +46% 0.19 ± 0.02 +64% walker-run 0.10 ± 0.01 0.32 ± 0.07 0.21 ± 0.05 +52% 0.18 ± 0.03 +36% humanoid-walk 0.02 ± 0.01 0.06 ± 0.01 0.05 ± 0.02 +67% 0.09 ± 0.02 +174% Mean DCS 0.13 ± 0.02 0.35 ± 0.04 0.24 ± 0.04 +50% 0.24 ± 0.03 +50% hammer 0.75 ± 0.07 0.98 ± 0.01 0.96 ± 0.01 +91% 0.99 ± 0.02 +102% bin-picking 0.18 ± 0.08 0.74 ± 0.10 0.49 ± 0.10 +56% 0.33 ± 0.08 +27% basketball 0.17 ± 0.03 0.51 ± 0.07 0.34 ± 0.09 +50% 0.37 ± 0.09 +58% soccer 0.14 ± 0.06 0.36 ± 0.08 0.21 ± 0.08 +34% 0.23 ± 0.06 +41% Mean DMW 0.31 ± 0.06 0.65 ± 0.06 0.50 ± 0.07 +55% 0.48 ± 0.06 +50%
Our experiments are designed to address the following core questions: (1) Can object-centric representations disentangle task-relevant motion from distracting visual elements? (2) How much supervision is needed to train effective latent action models under high visual distraction? (3) Can we automate slot selection to reduce the dependency on domain expertise or manual annotation?
4.1 Datasets
We evaluate our approach using two benchmark environments: one focused on control tasks under distractions (DCS) and another involving multi-object manipulation (DMW). More details on dataset collection can be found in Appendix A.
@@ 4.1 Datasets
The DCS environment (Stone et al., 2021) extends DeepMind Control with three types of real-world visual noise: (1) dynamic backgrounds: 60 diverse videos from DAVIS 2017 (Pont-Tuset et al., 2018), simulating realistic environmental clutter; (2) color variations: hue/saturation shifts that preserve object semantics but degrade low-level features; (3) camera perturbations: randomized pose adjustments mimicking handheld camera noise. We mainly focused on dynamic backgrounds (marked as DCS), but also showed the results on all 3 types of distraction (marked as DCS-hard) in Figures 1 and 7 and Table 1. We study four locomotion tasks ordered by complexity: cheetah-run, walker-run, hopper-hop, and humanoid-walk. We trained experts on privileged state information and collected the dataset with observation-action tuples. Importantly, the behavior cloning agent, trained on privileged true actions, is able to achieve expert score (Table 2).
Distracting MetaWorld (DMW).
The DMW environment extends MetaWorld with dynamic backgrounds, replicating the DCS setting. MetaWorld is a goal-based robotic multi-object benchmark featuring complex interactions between multiple objects. To our knowledge, MetaWorld with background distractions has not been previously explored. Tasks include hammer, bin-picking, basketball, and soccer. These tasks emphasize compositional reasoning and object interaction, allowing us to assess multi-object scenarios. The experts were taken from MetaWorlds’s oracle policies.
4.2 Setup
@@
To conduct the experiments, 4 models were trained: LAPO, LAPO-vanilla, LAPO-slots and LAPO-masks. All models were trained on the same datasets. More details on training can be found in Appendix B.
Baseline.
@@
The baseline model is LAPO, which is trained on observations with distractors, following the Schmidt & Jiang (2024) procedure (LAPO in the figures). We use it as a baseline to demonstrate the currently existing limitations of latent action pretraining.
Upper bound.
@@
We also trained LAPO on clean data without distractors as our second baseline to illustrate the performance gap caused by distractors (LAPO-vanilla in the figures).
Our approaches.
@@
LAPO-slots and LAPO-masks are the models that follow the object-centric latent action pre-training training pipeline described in Section 3 (respectively, LAPO-slots and LAPO-masks in the figures). LAPO-slots uses slot embeddings extracted from VideoSAUR as input to the latent action model, LAPO-masks uses slot masks, treating each mask as an attention map applied to the original frame.
4.3 Evaluation performance
To assess the quality of learned latent actions, we train a BC agent $\pi_{\text{BC}}(a_{t}\mid z_{t})$ on the inferred latent actions $z_{t}$, then fine-tune it on a small number of trajectories ($0.1\%-2.5\%$ from the entire dataset) with true action labels. The small size of the action-labeled sample mirrors real-world deployment, where labeling is expensive. The final performance is measured as normalized episodic return for DCS and normalized success rate for DMW, where a score of 1 corresponds to an oracle BC policy (see Table 2) trained on the entire dataset with ground-truth actions.
Quantitative Results.
@@ 4.3 Evaluation performance
As shown in Table 1, both LAPO-slots and LAPO-masks strongly outperform the standard LAPO baseline for both the DCS and DMW domains. Notably: In DCS, LAPO-slots improves the gap between LAPO performance with and without distractors by 50% on average, with even greater gains in complex tasks like ‘humanoid-walk‘ (+174%). In DMW, the improvements are consistent, with LAPO-slots achieving a +50% improvement on average over the baseline. These results support our hypothesis: object-centric representations help filter out distractors and retain actionable information. Figure 4 shows evaluation returns across increasing numbers of fine-tuning trajectories. Both slot-based methods show better generalization from fewer labeled examples.
4.4 Slot selection investigation
Figure 5: Slot selection investigation. From left to right: (1) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=4$. (2) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (3) Linear action probes and corresponding resulting normalized success rates of BC agents on 128 labeled trajectories for different number of slots (top-2 slots for $K\leq 5$, top-3 slots for $K\geq 8$).
A major benefit of object-centric modeling lies in its ability to attribute dynamics to specific entities. However, identifying which slots correspond to control-relevant entities remains a challenge. We explore whether linear probes trained to predict ground-truth actions from slot representations can identify relevant slots automatically. The detailed explanation of the technique is given in Section 3. This method shows strong agreement with manual selection and correlates well with downstream performance (see Figure 5 and Appendices E, F, G and H), suggesting that slot probing can serve as a reliable proxy for slot relevance.
Interestingly, we observe that meaningful task-related information tends to concentrate among the top-1 or top-2 slots (Figure 6). Specifically: for $K\leq 8$, only the top-1 slot contributes non-zero evaluation scores. For larger $K$ values ($K=11$ and $K=15$), the top-2 slots capture most of the useful signal. Aggregating the two most relevant slots through concatenation leads to improved performance (LABEL:fig:aggregated-performance), compared to using each slot individually. This suggests that meaningful information is distributed across these two slots.
Figure 6: Normalized success rates of the BC agent trained on latent actions for varying numbers of fine-tuning labeled trajectories across a range of slot counts from $K=2$ to $K=15$ on DMW tasks. Slots are ranked based on linear probe scores. TL;DR: Automatic probe selection correlates well with downstream performance. The values are averaged across three random seeds. The BC agent trained with access to the full dataset of ground-truth actions would return a score of 1 for each task.
4.5 Varying number of slots
The number of slots $K$ is a critical hyperparameter in object-centric models like VideoSAUR (Zadaianchuk et al., 2023), directly influencing both representation quality and downstream performance. Although prior work has explored adaptive slot allocation strategies (Fan et al., 2024), most object-centric approaches still rely on a fixed $K$, typically selected based on the expected complexity of the scene.
We evaluated LAPO-slots across a range of slot counts from $K=2$ to $K=15$ on DMW tasks. As shown in Figure 6, performance generally improves as $K$ increases, reaching its peak at 3–5 slots, after which it declines. Too few slots lead to collapsed representations in which multiple entities are merged into one, while too many introduce noise, redundancy, and instability (see Appendix H for visualizations). For instance, we observe "duplicating" slots (nearly identical representations) for $K\geq 5$, e.g., for $K=5$, only two unique slots emerge (for $K=8$ four distinct slots, for $K=11$ or $K=15$ only seven).
@@ 4.5 Varying number of slots
Interestingly, when applied to DMW, we expected a clear separation of distinct objects such as the robotic arm, hammer, ball, etc. Instead, VideoSAUR grouped these task-relevant entities into a single slot even when the number of slots $K$ was greater than the amount of objects. We attribute this behavior to limited diversity in expert trajectories: within each task, variation arises primarily from changes in the initial position of the manipulated object (e.g., hammer or ball), with no variation in the initial position of the arm or target object. This lack of data diversity likely limits the incentive for fine-grained decomposition and highlights a key limitation of current object-centric models: they struggle to disentangle objects meaningfully when the underlying data lack sufficient variability in object interactions or spatial configurations. It also shows that the setting of our study reflects real-world conditions, where data is rarely perfectly diverse or balanced.
4.6 Increasing distractions difficulty
@@
To better understand how different types of visual distractions impact latent action learning, we evaluate our methods under increasingly challenging conditions in the Distracting Control Suite (DCS) (Stone et al., 2021) environment. In this work, we primarily focus on dynamic backgrounds as the core challenge, but we also test combinations of all three distractor types (backgrounds, camera position, and color variations) on DCS (marked DCS-hard on Figure 1 and Table 1) to assess robustness.
As shown in Figure 1, all methods suffer performance degradation under additional distractors, for example, LAPO performance drops by x2.7 under background distractions and by x4.4 under additional distractors, compared to the performance without distractors. However, LAPO-slots maintains a consistent relative improvement over baseline (50% on DCS $\rightarrow$ 52% on DCS-hard), while LAPO-masks degradates (50% on DCS $\rightarrow$ 26% on DCS-hard).
@@
This difference probably stems from VideoSAUR’s use of the DINOv2 encoder Oquab et al. (2023), which captures high-level semantic features and is more robust to appearance changes than the CNN-based encoder in LAPO and LAPO-masks. These findings suggest that slot-based representations not only isolate task-relevant entities but also encode objects using features that generalize better under real-world visual perturbations.
@@ 5 Related work
Vision-Language-Action (VLA) models have emerged as a powerful paradigm for embodied AI, with recent advances focusing on latent action representations to enhance their capabilities. Ye et al. (2024) uses latent actions by pretraining VLM to predict latent actions from text prompt, enabling unsupervised pretraining of VLA models for robotic manipulation. Genie (Bruce et al., 2024) is a foundational world model, trained from unlabeled videos, which can generate interactive environments from images. Recently, Bu et al. (2025) trained cross-embodiment VLA policies for manipulation and navigation. They use LAM to derive task-centric action representations from videos.
Villar-Corrales & Behnke (2025) explores the application of object-centric representations for latent action learning and video prediction, showing positive results in robotic tabletop simulations. In contrast to our work, Villar-Corrales & Behnke (2025) does not explore the limitations of such an approach to latent action learning in the presence of distractors, similar to existing research. Moreover, instead of VideoSAUR (Zadaianchuk et al., 2023) employed in our work, their method uses SAVi (Kipf et al., 2021) for object-centric learning, which, as we show (see Figure 24) does not work well in more complex environments with real-world distractors.
Nikulin et al. (2025) focuses on latent action learning under distractors, closely aligned with ours. Even though the topic is the same, the method, presented by the authors is ortogonal. It presents a latent action framework without reconstruction which shows improvement with distractors, however, utilizes supervision (small dataset of labeled trajectories) during pre-training.
@@
Together, these works highlight the growing interest in robust latent action learning under visual distractions. Our contribution uniquely bridges object-centric perception and latent action modeling, emphasizing slot stability, interpretability, and reduced supervision requirements.
6 Conclusion & Limitations
@@
Our results demonstrate that object-centric representations significantly mitigate the impact of distractors in learning latent actions from videos. By disentangling scene elements into meaningful, interpretable slots, latent actions focus on causal dynamics rather than spurious correlations. This could provide a strong inductive bias for more effective Latent Action Models in noisy environments. However, several limitations remain that constrain the applicability and scalability of current object-centric frameworks.
Despite being unsupervised, object-centric models like VideoSAUR do not offer direct control over what is learned. Although we successfully automate slot selection using linear probing, this process still requires access to a small set of labeled trajectories to train the probe itself. Similarly, while latent action learning benefits from large-scale unlabeled video data, fine-tuning on expert-labeled trajectories remains essential for achieving strong downstream performance. Future work may explore other directions for incorporating weak supervision.
A key limitation we observe is the sensitivity of object-centric models to dataset composition. In robotic manipulation tasks, such as DMW, the limited diversity in expert trajectories leads to collapsed object representations. A promising direction towards mitigating data diversity issue is the use of generative approaches (Luo et al., 2025; Sikchi et al., 2024) to augment the dataset, enabling models to learn more robust and generalizable object decompositions through knowledge transfer.
A major limitation of current slot-based video models like VideoSAUR is the lack of memory mechanisms, which restricts modeling dynamic scenes where objects may enter or leave the scene over time. This issue becomes especially important in long-horizon sequences involving occlusions, multi-object interactions, or complex environments. Recent attempts to address this include dictionary-based slot architectures and hierarchical slot structures (Đukić et al., ), but these remain early-stage and mostly limited to image-level processing rather than full video understanding. Furthermore, most OCL methods operate under the assumption of a known and fixed number of objects $K$, limiting their ability to model hierarchical structures or scenes with varying numbers of objects.
[…]

### v3

@@ 4 Experiments
Our experiments are designed to address the following core questions: Can object-centric representations disentangle task-relevant motion from distracting visual elements? How much supervision is needed to train effective latent action models under high visual distraction? Can we automate slot selection to reduce the dependency on domain expertise or manual annotation?
Datasets
We evaluate on two benchmark environments: Distracting Control Suite (DCS) and Distracting Meta-World (DMW). Details on data collection are in Supplementary Material.
@@
DCS (Stone et al. 2021) extends DeepMind Control with three visual distractions: (1) dynamic backgrounds (DAVIS 2017 videos), (2) color variations (hue/saturation shifts), and (3) camera perturbations. We focus on dynamic backgrounds (DCS) and report full results on all distractions (DCS-hard) in Figures˜1 and 1 and Supplementary Material. Tasks: cheetah-run, walker-run, hopper-hop, humanoid-walk (increasing complexity). Experts were trained on privileged state; datasets consist of observation-action pairs. Behavior cloning, trained on privileged true actions, is able to achieve expert performance.
Distracting Meta-World (DMW).
DMW extends Meta-World (Yu et al. 2019) with DCS-style dynamic backgrounds. Tasks (hammer, bin-picking, basketball, soccer) involve multi-object interactions and compositional reasoning. Experts use Meta-World’s oracle policies.
Figure 5: Slot selection study on basketball task. (a,b) Linear action probes under varying labeled trajectories budget and corresponding examples of slot masks for basketball task for $K=4$. (c) BC performance on different slots for $K=4$ and $K=15$ on basketball task. (d) Linear action probe scores vs. normalized BC success rates (on 128 trajectories). TL;DR: Probe-based selection correlates with downstream performance. (e) BC performance on varying $K$ parameter on basketball task. Full-action BC achieves score of 1. TL;DR: Regardless of the $K$ parameter LAPO-slots outperform baseline LAPO.
Models
@@
To conduct the experiments, 4 models were trained: LAPO, LAPO-clean, LAPO-slots and LAPO-masks. All models were trained on the same datasets. More details on training can be found in Supplementary Material.
@@
The baseline model is LAPO, which is trained on observations with distractors, following the Schmidt and Jiang (2024) procedure (LAPO in the figures). We use it as a baseline to demonstrate the currently existing limitations of latent action pretraining.
@@
We also trained LAPO on clean data without distractors as our upper bound to illustrate the performance gap caused by distractors (LAPO-clean in the figures).
@@
LAPO-slots and LAPO-masks are the models that follow the object-centric latent action pre-training pipeline described in Section˜3 (respectively, LAPO-slots and LAPO-masks in the figures). LAPO-slots uses slot embeddings extracted from VideoSAUR as input to the latent action model, LAPO-masks uses slot masks, treating each mask as an attention map applied to the original frame.
Evaluation performance
@@
To assess the quality of learned latent actions, we train a BC agent $\pi_{\text{BC}}(a_{t}\mid z_{t})$ on the inferred latent actions $z_{t}$, then fine-tune it on a small number of trajectories ($0.1\%-2.5\%$ from the entire dataset) with true action labels. The small size of the action-labeled sample mirrors real-world deployment, where labeling is expensive. The final performance is measured as normalized episodic return for DCS and normalized success rate for DMW, where a score of 1 corresponds to an oracle BC policy (see Supplementary Material) trained on the entire dataset with ground-truth actions. As shown in Table˜1, both LAPO-slots and LAPO-masks strongly outperform the standard LAPO baseline for both the DCS and DMW domains. Notably: In DCS, LAPO-slots improves the gap between LAPO performance with and without distractors by 50% on average, with even greater gains in complex tasks like ‘humanoid-walk‘ (+174%). In DMW, the improvements are consistent, with LAPO-slots achieving a +50% improvement on average over the baseline. These results support our hypothesis: object-centric representations help filter out distractors and retain actionable information. Figure˜4 shows evaluation returns across increasing numbers of fine-tuning trajectories. Both slot-based methods show better generalization from fewer labeled examples.
Slot Selection Investigation
Object-centric models enable interpretable, modular control by attributing dynamics to individual entities, but identifying which slots encode control-relevant information remains challenging.
We show that linear action probes, trained to predict actions from individual slot representations, can automatically identify relevant slots (method in Section˜3). Probe scores align with manual analysis (Figures˜5(a) and 5(b)) and strongly correlate with downstream BC performance (Figure˜5(d)), validating them as a reliable relevance metric.
Probing reveals that useful information concentrates in either a small set of dominant slots (one slot for $K\leq 8$ or a tightly coupled pair for $K>8$), with sharply worse scores for all others. Only these top-ranked slots yield non-zero BC performance (Figure˜5(c) for $K=4$ and $K=15$ cases); the rest are ineffective. For $K>8$, concatenating the top two slots improves performance, confirming they capture complementary information. Thus, probe score distributions guide both selection and fusion of meaningful slots.
Additional visualizations are presented in the Supplementary Material.
Varying Number of Slots
The number of slots $K$ is a key hyperparameter in object-centric models like VideoSAUR (Zadaianchuk et al. 2023), often set based on scene complexity. While adaptive allocation has been explored (Fan et al. 2024), most methods use a fixed $K$, making robustness to its choice crucial for real-world deployment. We evaluate LAPO-slots for $K=2$ to $15$ on DMW tasks. As shown in Figure˜5(e), performance remains consistently strong across all $K$, demonstrating robustness to the choice of $K$, which is a practical advantage when the number of relevant entities is unknown. Too few slots ($K=2$) cause collapsed representations, merging multiple entities into one. In contrast, larger $K$ introduces duplicated slots with nearly identical representations. For $K=5$, only two unique slots emerge; for $K=8$, four; for $K=11$ or $15$, seven. Despite redundancy, downstream performance remains stable, as slot selection focuses on the most informative ones.
@@
Interestingly, when applied to DMW, we expected a clear separation of distinct objects such as the robotic arm, hammer, ball, etc. Instead, VideoSAUR grouped these task-relevant entities into a single slot even when the number of slots $K$ was greater than the amount of objects. We attribute this to limited trajectory diversity: variations arise from the manipulated object’s initial position, with little change in arm or target configuration. Without sufficient variability, the model lacks incentive to disentangle co-moving entities. This reflects a general tendency in object-centric models: they favor functional over geometric decomposition, merging semantically distinct but behaviorally coupled entities. Importantly, this behavior mirrors real-world conditions, where data is often imbalanced and incomplete.
Increasing distractions difficulty
@@
To better understand how different types of visual distractions impact latent action learning, we evaluate our methods under increasingly challenging conditions in the Distracting Control Suite (DCS) (Stone et al. 2021) environment. In this work, we primarily focus on dynamic backgrounds as the core challenge, but we also test combinations of all three distractor types (backgrounds, camera position, and color variations) on DCS (marked DCS-hard on Figure˜1 and Table˜1) to assess robustness.
@@
As shown in Figure˜1, all methods suffer performance degradation under additional distractors, for example, LAPO performance drops by x2.7 under background distractions and by x4.4 under additional distractors, compared to the performance without distractors. However, LAPO-slots maintains a consistent relative improvement over baseline (50% on DCS $\rightarrow$ 52% on DCS-hard), while LAPO-masks degrades (50% on DCS $\rightarrow$ 26% on DCS-hard). This difference probably stems from VideoSAUR’s use of the DINOv2 encoder (Oquab et al. 2023), which captures high-level semantic features and is more robust to appearance changes than the CNN-based encoder in LAPO and LAPO-masks. These findings suggest that slot-based representations not only isolate task-relevant entities but also encode objects using features that generalize better under real-world visual perturbations.
@@ 5 Related work
Vision-Language-Action (VLA) models like LAPA (Ye et al. 2024), Genie (Bruce et al. 2024), and UniVLA (Bu et al. 2025) have shown impressive capabilities by training on large-scale video data and deploying on real robots. However, scale can make it difficult to isolate specific failure modes. Our work focuses on one such critical failure: the sensitivity of the underlying Latent Action Models to visual distractors. To enable a controlled study of this problem, we use the Latent Action Policy Optimization (LAPO) framework (Schmidt and Jiang 2024) within simulated environments. While effective in clean settings, LAPO’s performance degradation under distraction makes it an ideal baseline to rigorously evaluate robustness improvements.
Our work confronts this distractor challenge by integrating object-centric learning (OCL). OCL has shown significant promise for compositional generation (Akan and Yemez 2025), building structured world models (Heins et al. 2025; Barcellona et al. 2024) and compositional video prediction (Lee et al. 2025). The application of OCL to latent action learning has been explored by Villar-Corrales and Behnke (2025), who use object-centric models for latent action learning in robotic simulations. However, their work, along with most OCL-based world models, does not investigate performance under heavy visual distraction, a key focus of our paper. Furthermore, their reliance on the SAVi architecture (Kipf et al. 2021) for object discovery proves less robust in complex, distracting environments than the VideoSAUR model (Zadaianchuk et al. 2023) we employ (see Supplementary Material).
@@
Another work that directly addresses latent action learning with distractors is Nikulin et al. (2025). While their goal is similar, their method is orthogonal: they propose a reconstruction-free framework that requires a small set of labeled trajectories to provide supervision during pre-training. Together, these works highlight the growing interest in robust latent action learning under visual distractions. Our contribution uniquely bridges object-centric perception and latent action modeling, emphasizing slot stability, interpretability, and reduced supervision requirements.
6 Discussion and Limitations
@@
Our results demonstrate that object-centric (OC) representations can significantly mitigate the impact of distractors when learning latent actions from video. By disentangling scenes into meaningful slots, our approach allows latent action models to focus on causal dynamics rather than spurious correlations, providing a strong inductive bias for learning in noisy environments.
The effectiveness of our method is inherently linked to the capabilities of the underlying OCL model. We believe that as OCL methods continue to improve and scale from simulated to more complex, real-world data, much as VideoSAUR improved upon previous models, so too will the performance and applicability of our approach. Our contribution is to show that integrating object decomposition is a viable path for robust latent action learning, thereby motivating further investment in the underlying OCL technologies.
The reliance on current OCL frameworks introduces several limitations. While powerful segmentation models could isolate objects, they are not inherently designed to handle the dynamic, action-correlated distractors (e.g., camera motion, color shifts) which are explored in our work. The challenges become even greater in scenarios with heavy occlusions or multiple camera viewpoints. Furthermore, current slot-based models like VideoSAUR lack memory mechanisms to handle objects entering or leaving the scene, and often assume a fixed number of objects K. While recent work has explored solutions like dictionary-based architectures and adaptive slot allocation (Đukić et al. 2025; Fan et al. 2024), these remain open challenges.
Finally, we observed that the unsupervised nature of OCL models, particularly under limited data diversity, can lead to non-intuitive collapsed decompositions due to lack of direct control over the learned representations. A promising direction to mitigate this is to augment training data with generative approaches (Luo et al. 2025; Sikchi et al. 2024) or develop weakly-supervised OCL methods to learn more controllable representations.
Acknowledgments
This work was supported by the The Ministry of Economic Development of the Russian Federation in accordance with the subsidy agreement (agreement identifier  [^00] C313925P4H0002; grant No 139-15-2025-012).
Supplementary Material